# Enhancing Textual Personality Detection toward Social Media: Integrating Long-term and Short-term Perspectives

### v1

@@ 4.3 Implementation Details
For training, we employed the Adam optimizer [15] with a learning rate of 2e-5 for the pretrained BERT in the Short-term Personality Encoding layer and 2e-3 for the other components of the models. Mini-batch size was set to 32. The training process continued for a maximum of 10 epochs, and the model exhibiting the best performance on the validation set was chosen for testing. To maintain result consistency and minimize experimental variations, we set all random seeds used in the experiments to 1.
Approch Kaggle Pandora
E/I S/N T/F J/P Average E/I S/N T/F J/P Average
SVM 53.34 47.75 76.72 63.03 60.21 44.74 46.92 64.62 56.32 53.15
XGBoost 56.67 52.85 75.42 65.94 62.72 45.99 48.93 63.51 55.55 53.50
BiLSTM 57.82 57.87 69.97 57.01 60.67 48.01 52.01 63.48 56.12 54.91
AttRCNN 59.74 64.08 78.77 66.44 67.25 48.55 56.19 64.39 57.26 56.60
BERT 65.02 59.56 78.45 65.54 67.14 56.60 48.71 64.70 56.07 56.52
SN + Attn 62.34 57.08 69.26 63.09 62.94 54.60 49.19 61.82 53.64 54.81
Trans-MD 66.08 69.10 79.19 67.50 70.47 55.26 58.77 69.26 60.90 61.05
TrigNet 69.54 67.17 79.06 67.69 70.86 56.69 55.57 66.38 57.27 58.98
D-DGCN 67.37 64.47 80.51 66.10 69.61 58.28 55.88 68.50 57.72 60.10
DEN(Ours) 69.95 66.39 80.65 69.02 71.50 60.86 57.74 71.64 59.17 62.35
Table 1: Overall Macro-F1 (%) results of DEN and baseline Models on the Kaggle dataset and Pandora dataset, where the best results are shown in bold, and the second best are underlined.
5 Overall Results
The overall results for DEN and the baseline models are presented in Table 1. These results demonstrate that our proposed DEN outperforms other models, achieving the highest average F1 score in both the Kaggle and Pandora datasets. Particularly noteworthy is the relative improvement of 2% in the Pandora dataset when compared to the previous state-of-the-art models. This underscores the effectiveness of our DEN in personality detection.
@@ 5 Overall Results
6 Detailed Analysis
The overall results presented above have demonstrated the effectiveness of our DEN model as a whole. To gain deeper insights and understand the impact of each key module in DEN, we conducted a series of experiment on various variants of DEN, and the general results are shown in Table 2. It can be observed that removing any part of the model leads to a performance decrease in DEN. The detailed findings and analysis will be discussed in the following sections.
@@ 6 Detailed Analysis
Table 2: Results of various variants of DEN model in Macro-F1(%) on the test set of Kaggle dataset, where “–” denotes the removal of a component.
6.1 Effect of Long-term Personality Encoding
In the variant of DEN, denoted as DEN-long, we removed the entire Long-term Personality Encoding layer, resulting in a 4.36% decrease in performance. This indicates the crucial importance of considering long-term information in personality detection. In the variant DEN-gcn, we removed the graph convolutional neural network module in the Long-term Personality Encoding layer, which led to a 2.4% decrease in performance. This emphasizes the significance of modeling long-term user information for personality detection. These provide evidence that our DEN effectively utilizes graph convolutional neural networks to analyze patterns in users’ usage of psychological entities, allowing it to efficiently acquire users’ long-term personality representations and improve personality detection performance.
@@ 6.1 Effect of Long-term Personality Encoding
6.2 Effect of Short-term Personality Encoding
@@
When the short-term Personality Encoding module is removed, denoted as DEN-short, the performance drops dramatically, almost to an unusable level. This is because, without the short-term Personality Encoding module, the model only considers users’ habits in using psychological entities, leading to a significant loss of semantic information. While these habits provide some level of user modeling, the absence of semantic information makes it difficult to predict user personalities. Furthermore, the model is hindered by a scarcity of parameters at this point, which poses challenges in accurately fitting the data. Therefore, similar to previous work, DEN utilizes pre-trained language models and the user’s textual context information to model the user’s short-term representations, enriching user information and thus enhancing personality detection performance.
6.3 Effect of Bi-directioan Interaction
@@
Figure 7: Effect of Bi-directioan Interaction
@@
6.4 Effect of External Knowlege Base
@@
Figure 8: Effect of External Knowlege Base
6.5 Effect of Gated Fusion Mechanism
@@
7 Conclusion and Future Work
@@
In our current research, both our psychological theoretical analysis and the DEN personality detection model were based on the MBTI personality model. However, we recognize the importance of exploring the effectiveness of the DEN model on other personality models, such as the Big Five, in future work. By examining the applicability of the DEN model to different personality frameworks, we can broaden our understanding of its performance and its ability to generalize across various personality models.

### v2

@@ IV-C Implementation Details
For training, we employed the Adam optimizer [39] with a learning rate of 2e-5 for the pre-trained BERT in the Short-term Personality Encoding layer and 2e-3 for the other components of the models. Mini-batch size was set to 32. The training process continued for a maximum of 10 epochs, and the model exhibiting the best performance on the validation set was chosen for testing. To maintain result consistency and minimize experimental variations, we set all random seeds used in the experiments to 1.
Approach E/I S/N T/F J/P Average
SVM 53.34 47.75 76.72 63.03 60.21
XGBoost 56.67 52.85 75.42 65.94 62.72
BiLSTM 57.82 57.87 69.97 57.01 60.67
AttRCNN 59.74 64.08 78.77 66.44 67.25
BERT 65.02 59.56 78.45 65.54 67.14
SN + Attn 62.34 57.08 69.26 63.09 62.94
Trans-MD 66.08 69.10 79.19 67.50 70.47
TrigNet 69.54 67.17 79.06 67.69 70.86
D-DGCN 67.37 64.47 80.51 66.10 69.61
DeepSeek V3 63.31 54.88 74.20 50.48 60.72
GPT-4o 64.59 60.74 74.92 50.18 62.61
PsyCoT 66.56 61.70 74.80 57.83 65.22
DEN(Ours) 69.95* 66.39 80.65* 69.02* 71.50*
TABLE I: Overall Macro-F1(%) results of DEN and baseline models on the Kaggle dataset, where the best results are shown in bold, and the second-best are underlined. Significant improvements over the best baseline are marked with * ($p<0.05$), as determined by a paired t-test.
V Overall Results
The overall results for DEN and the baseline models are presented in Tables I and II. These results demonstrate that our proposed DEN outperforms other models, achieving the highest average F1 score on both the Kaggle and Pandora datasets. Notably, DEN achieves a relative improvement of 2% on the Pandora dataset compared to the previous state-of-the-art models, underscoring its effectiveness in personality detection.
@@ V Overall Results
Approach E/I S/N T/F J/P Average
SVM 44.74 46.92 64.62 56.32 53.15
XGBoost 45.99 48.93 63.51 55.55 53.50
BiLSTM 48.01 52.01 63.48 56.12 54.91
AttRCNN 48.55 56.19 64.39 57.26 56.60
BERT 56.60 48.71 64.70 56.07 56.52
SN + Attn 54.60 49.19 61.82 53.64 54.81
Trans-MD 55.26 58.77 69.26 60.90 61.05
TrigNet 56.69 55.57 66.38 57.27 58.98
D-DGCN 58.28 55.88 68.50 57.72 60.10
DEN(Ours) 60.86* 57.74 71.64* 59.17 62.35*
TABLE II: Overall Macro-F1(%) results of DEN and baseline models on the Pandora dataset.
@@
However, we also observed that all models, including DEN, perform relatively poorly on the Sensing/Intuition (S/N) personality dimension. Moreover, DEN’s advantage in this dimension is less pronounced compared to other dimensions. We hypothesize that this is due to the highly imbalanced nature of personality detection datasets, with the S/N dimension being particularly skewed. As DEN integrates both long-term and short-term modeling, it may have inadvertently amplified these biases. To address this issue, future work could explore strategies such as data augmentation, class rebalancing, or the use of focal loss to mitigate the impact of data imbalance.
VI Detailed Analysis
The overall results presented above have demonstrated the effectiveness of our DEN model as a whole. To gain deeper insights and understand the impact of each key module in DEN, we conducted a series of experiment on various variants of DEN, and the general results are shown in Table III. It can be observed that removing any part of the model leads to a performance decrease in DEN. The detailed findings and analysis will be discussed in the following sections.
@@ VI Detailed Analysis
TABLE III: Results of variants of DEN model in Macro-F1(%) on the test set of Kaggle dataset.
VI-A Effect of Long-term Personality Encoding
@@
In the variant DEN-long, we removed the entire Long-term Personality Encoding layer and no long-term information was incorporated into the modeling process, resulting in a 4.36% decrease in performance. This indicates the crucial importance of considering long-term information in personality detection. In the variant DEN-gcn, we removed the graph convolutional neural network module in the Long-term Personality Encoding layer but still fed the raw psychological entity representations into the Bi-directional Interaction module, which led to a 2.4% decrease in performance. This emphasizes the significance of modeling long-term user information for personality detection. These provide evidence that our DEN effectively utilizes graph convolutional neural networks to analyze patterns in users’ usage of psychological entities, allowing it to efficiently acquire users’ long-term personality representations and improve personality detection performance.
VI-B Effect of Short-term Personality Encoding
@@
When the short-term Personality Encoding module is removed, denoted as DEN-short, the performance drops dramatically, almost to an unusable level. This is because, without the short-term Personality Encoding module, the model only considers users’ habits in using psychological entities, leading to a significant loss of semantic information. While these habits provide some level of user modeling, the absence of post-level semantic summary vectors makes it difficult to predict user personalities. Furthermore, the model is hindered by a scarcity of parameters at this point, which poses challenges in accurately fitting the data. Therefore, similar to previous work, DEN utilizes pre-trained language models and the user’s textual context information to model the user’s short-term representations, enriching user information and thus enhancing personality detection performance.
VI-C Effect of Bi-directional Interaction
@@
Figure 7: Effect of Bi-directional Interaction
@@
VI-D Effect of External Knowledge Base
@@
Figure 8: Effect of External Knowledge Base
Approach INFP INFJ INTP INTJ ISFP ISFJ ISTP ISTJ ENFP ENFJ ENTP ENTJ ESFP ESFJ ESTP ESTJ
Support (%) 21.27 17.18 15.68 11.12 3.52 2.19 4.15 2.07 8.01 1.84 8.18 2.42 0.52 0.52 1.04 0.29
D-DGCN 53.69 41.09 46.79 34.43 25.58 15.39 33.33 15.39 32.23 9.88 39.73 25.32 0.00 0.00 6.67 0.00
DEN(Ours) 54.18 45.89 52.03 37.46 19.97 16.14 29.79 22.95 34.10 11.59 38.69 26.54 0.00 0.00 12.35 0.00
TABLE IV: Results of DEN and D-DGCN in Macro-F1(%) on the test set of Kaggle dataset on various personality combinations. Support refers to the ratio of support sets for each personality combination.
VI-E Effect of Gated Fusion Mechanism
@@
VI-F Effect of Number of Layers of Bi-Directional Interaction
Figure 10: Effect of Number of Layers of Bi-Directional Interaction
The short-term and long-term features in DEN exist in distinct spaces, and increasing the number of interaction layers may help integrate these features more effectively by reducing the gap between these spaces. To investigate the effect of the number of layers in the Bi-Directional Interaction component, we conducted a series of experiments, with the results presented in Figure 10. The results show that the model achieves its best performance with two interaction layers. Beyond this point, performance declines as additional layers are added. This improvement with two layers can be attributed to the model’s ability to capture higher-order neighborhood information, enhancing the integration of long-term and short-term features. However, adding too many layers results in over-smoothing, where node representations become indistinguishable, ultimately degrading the model’s performance. Considering these trade-offs, we adopt a two-layer network as it balances effective feature integration with avoiding over-smoothing, ensuring optimal model performance.
VI-G Performance on Personality Combinations
The MBTI personality framework consists of 16 unique personality combinations. To analyze the performance of DEN in greater detail, we report the results of DEN and D-DGCN across these combinations, as shown in Table IV. Our findings indicate that DEN outperforms D-DGCN in 13 out of 16 combinations, demonstrating the superiority of DEN in personality detection. This highlights the effectiveness of modeling personality through both long-term and short-term aspects.
However, the results also reveal a significant limitation: both DEN and D-DGCN perform poorly in personality combinations with limited data. In combinations representing less than 1% of the dataset, the F1-score for both models drops to 0, indicating a complete inability to make accurate predictions. These findings underscore the need for further research to address data scarcity in underrepresented personality combinations.
VII Conclusion and Future Work
@@
In our current research, both our psychological theoretical analysis and the DEN personality detection model were based on the MBTI personality model. However, we recognize the importance of exploring the effectiveness of the DEN model on other personality models, such as the Big Five, in future work. By examining the applicability of the DEN model to different personality frameworks, we can broaden our understanding of its performance and its ability to generalize across various personality models. Another important direction for future research is addressing the challenge of imbalanced data in personality detection tasks. Many personality traits and combinations are underrepresented in existing datasets, which can lead to biased model performance and poor generalization, particularly for less frequent classes. While our study employed Macro-F1 as an evaluation metric to mitigate the influence of imbalance, this does not fully resolve the issue. Future work could explore advanced methods, such as data augmentation, re-sampling strategies, or loss functions like focal loss, to better address class imbalance and improve the robustness of personality detection models across all personality categories. These enhancements would contribute to more equitable and accurate personality prediction, even for underrepresented traits and combinations.
VIII Acknowledgements
This research is supported by the National Natural Science Foundation of China (No. 62376051).